a=300, h-25
a-bh/2
trying to find base

Answer by checkley71(8403)

IF YOU MEAN AREA=BASE*H/2 THEN 300=B*25/2 OR B=300/(2/25) OR B=300*2/25 OR
B=600/25 OR B=24
PROOF 300=24*25/2 OR 300=600/2 OR 300=300